Circular No. 7297

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999en, 1999eo, 1999ep, 1999eq, 1999er L.-G. Strolger, University of Michigan and Cerro Tololo Interamerican

Observatory (CTIO); R. C. Smith, CTIO; E. P. Rubenstein, CTIO; A. Soderberg, Bates College and CTIO; and T. Bowers, University of Arizona, report the discovery by the Nearby Galaxies Supernova Search team of the following supernovae in their Sept. 27–Oct. 4 search using the Kitt Peak 0.9-m telescope (+ Mosaic imager):

SN 1999 UT α_{2000} δ_{2000} R Offset
1999en Oct. 3 1^{h}15^{m}29.^{s}70 − 4^{o}30′48′′.2 19 : 0′′.5 E, 4′′.2 N
1999eo Oct. 2 2 40 13.20 + 4 54 53.3 19.6 4′′.4 W, 3′′.0 S
1999ep Sept. 28 4 41 04.76 − 3 00 39.6 18.9 6′′.5 W, 3′′.4 S
1999eq Oct. 3 1 19 22.70 − 3 36 19.8 19 : 5′′.4 E, 10′′.5 N
1999er Oct. 1 2 47 06.88 − 2 57 43.9 20.4 3′′.0 E, 37′′.7 N

SN 1999en is identified as a type-Ia event from spectra (range 400–800 nm) obtained by Strolger and Smith with the 4-m Mayall telescope on Oct. 8–10; SN 1999en was ∼ 7 days past maximum light on Oct. 16 in spectra obtained by M. M. Phillips with the 2.5-m Dupont telescope on Oct. 15–16. SN 1999en has a redshift of ∼ 0.16. Spectra obtained on the same dates at Kitt Peak and Las Campanas show that SN 1999eo is a type-Ia supernova, and > 30 days past maximum light on Oct. 16, with a redshift of ∼ 0.08. SN 1999ep is a type-Ia event, ∼ 7 days past maximum light on Oct. 8, at a redshift of ∼ 0.06, as identified from spectra obtained at Kitt Peak. SN 1999eq is a type-II event based on spectra (range 400–800 nm) obtained by Strolger and Smith with the 4-m Mayall telescope on Kitt Peak on Oct. 8–10 and by Phillips using the Dupont 2.5-m telescope at Las Campanas on Oct. 15–16; SN 1999eq (redshift ∼ 0.05) exhibits a strong Hα emission peak near 690 nm, preceded by a relatively weak absorption trough. From spectra obtained on the same dates at Kitt Peak and Las Campanas, SN 1999er in MCG –01-08-008 is identified as a young type-II supernova, < 14 days old on Oct. 1. SN 1999er has well-defined hydrogen P-Cyg profiles with Hα emission peaking ∼ 670 nm. Photometric confirmation of these supernovae were made by A. Athey and Strolger using the Michigan-Dartmouth-M.I.T. 2.4-m telescope on Oct. 4–7, as well as by P. Candia and R. Covarrubias using the CTIO 0.9-m telescope on Oct. 5 and 7. R magnitudes on Oct. 20: SN 1999en, 20.5; SN 1999eo, 19.8; SN 1999ep, 19.5; SN 1999eq, 20.3; SN 1999er, 19.8. Finding charts and further information are located at

http://ctios6.ctio.noao.edu/~ngss/.

1999 November 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green